OpenStudy (anonymous):

FIND SLOPE (-2, 1) and ( 1/4 , 5)

OpenStudy (amistre64):

move from one point to the other :)

OpenStudy (amistre64):

slope is how y moves, divided by, how x moves

OpenStudy (amistre64):

y moves from 1 to 5; thats +4 x moves from -2 to 1/4; that +2' 1/4

OpenStudy (amistre64):

slope = 4/2' 1/4 = 4/ 9/4 = 16/9 maybe?

OpenStudy (anonymous):

(5-1)/(1/4+2) = 4/(9/4) =4*4/9 = 16/9

OpenStudy (anonymous):

thank you!!! :)

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

its ok